one factor of 18x^2-29x-40 is 2x-5. what is the other factor?

Answer by checkley79(3341)

18x^2-29x-40 is 2x-5.
18/2=9
-40/-5=8
THUS THE 2 MISSING NUMBERS.
OR:
(9X+8)
PROOF:
2X-5
9X+8 MULTIPLY
-----------------
18X^2-45X+16X-40
18X^2-29X-40